OpenStudy (anonymous):

(cos(theta)+1/square root of 2 )(cos(theta)+1/square root of 2 )=0

OpenStudy (experimentx):

solve for \( \theta\) ??

OpenStudy (anonymous):

yeah :)

OpenStudy (anonymous):

\[(\cos \theta+1/\sqrt{2})(\cos \theta+1/\sqrt{2})=0\]

OpenStudy (campbell_st):

\[\cos(\theta) = \pm \frac{1}{\sqrt{2}}\] then \[\theta = 45, 135, 225, 315\]

OpenStudy (anonymous):

ohh.. that's all ?

OpenStudy (experimentx):

either, \( \cos \theta + 1/\sqrt2 = 0\) or \( \cos \theta + 1/\sqrt2 = 0\) \( \cos \theta + \cos (2n\pi + \pi/4) = 0\)

OpenStudy (anonymous):